# Optimal Multi-Distribution Learning

### v2

@@ C.2.2 Extracting regular segments from irregular segments
For $11\leq x\leq\log_{2}(T)+$ 1, we define $\mathcal{W}_{j}(x):=\{i|2^{x-1}\leq e_{i}-s_{i}\leq 2^{x}\}$ and let $x^{*}=\arg\max_{1\leq x\leq\log_{2}(T)+1}|\mathcal{W}_{j}(x^{*})|$. Then we have that $|\mathcal{W}_{j}(x^{*})|\geq\frac{|\mathcal{W}_{j}|}{\log_{2}(T)+1}$. Then we consider to choose the subsets $\{\mathcal{W}_{j}^{m}\}_{m=1}^{M}$ from $\mathcal{W}_{j}(x^{*})$.
Let $\kappa_{1}=\min_{i\in\mathcal{W}_{j}(x^{*})}e_{i}$. We also define $\mathcal{U}_{j}^{1}:=\{i\in\mathcal{W}_{j}(x^{*})|s_{i}\leq\kappa_{1}\}$. For $o\geq 1$, we let
$\kappa_{o+1}=\min_{i\in\mathcal{W}_{j}(x^{*})/\cup_{o^{\prime}=1}^{o}\mathcal{U}_{j}^{o^{\prime}}}e_{i}$
$\mathcal{U}_{j}^{o+1}:=\{i\in\mathcal{W}_{j}(x^{*})/\cup_{o^{\prime}=1}^{o}\mathcal{U}_{j}^{o^{\prime}}|s_{i}\leq\kappa_{o+1}\}.$
And we stop until $\cup_{o\geq 1}\mathcal{U}_{j}^{o}=\mathcal{W}_{j}(x^{*})$.
Clearly, for each $o$ and $i_{1},i_{2}\in\mathcal{U}_{j}^{o}$, we have that $e_{i_{1}}\geq\kappa_{o}\geq s_{i_{2}}$. That is, $\max_{i\in\mathcal{U}_{j}^{o}}s_{i}\leq\min_{i\in\mathcal{U}_{j}^{o}}e_{i}$. We then consider the three groups of subsets of $\mathcal{W}_{j}(x^{*})$: $\{\mathcal{U}_{j}^{3o-2}\}_{o\geq 1}$, $\{\mathcal{U}_{j}^{3o-1}\}_{o\geq 1}$ and $\{\mathcal{U}_{j}^{3o}\}_{o\geq 1}$. Then there exists $i\in\{0,1,2\}$ such that $\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-i}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}$. Without loss of generality, we assume that $\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-2}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}=|\mathcal{W}_{j}(x^{*})|$.
Now we verify that $\{\mathcal{U}_{j}^{3o-2}\}_{o\geq 1}$ is the desired group of sets. The first condition holds because of the definition of $\{\mathcal{U}_{j}^{o}\}_{o\geq 1}$. As for the second condition, we have that
$\sum_{o\geq 1}|\mathcal{U}_{j}^{3o-2}|\geq\frac{1}{3}\sum_{o\geq 1}\mathcal{U}_{j}^{o}=\frac{1}{3}|\mathcal{W}_{j}(x^{*})|\geq\frac{|\mathcal{W}_{j}|}{3(\log_{2}(T)+1)}.$
As for the last condition, it suffices to verify that $\max_{i\in\mathcal{U}_{j}^{3o-2}}e_{i}\leq\min_{i\in\mathcal{U}_{j}^{3o+1}}s_{i}$ for any proper $o$. By definition, for each proper $o\geq 1$, there exists $i\in\mathcal{W}_{j}(x^{*})$ such that $s_{i}\geq\kappa_{o}$ and $\kappa_{o+1}=e_{i}$. Then we have that $\kappa_{o+1}=e_{i}\geq s_{i}+2^{x^{*}-1}\geq\kappa_{o}+2^{x^{*}-1}$. It then follows that, for any $i\in\mathcal{U}_{j}^{3o+1}$, $s_{i}\geq\kappa_{3o}\geq\kappa_{3o-2}+2^{x^{*}}$. On the other hand, for any $i\in\mathcal{U}_{j}^{3o-2}$, we have that $e_{i}\leq s_{i}+2^{x^{*}}\leq\kappa_{3o-2}+2^{x^{*}}$. Putting all together, we have that
$\max_{i\in\mathcal{U}_{j}^{3o-2}}e_{i}\leq\kappa_{3o-2}+2^{x^{*}}\leq\min_{i\in\mathcal{U}_{j}^{3o+1}}s_{i}.$
The proof is finished.
@@
Recall the definition of $\mathcal{W}_{j}$ and $(s_{i},e_{i})$ for $i\in\mathcal{W}_{j}$. Then there exists a group of subsets $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ such that the conditions below hold
  1. (i).
    $\mathcal{V}_{j}^{n}\subset\mathcal{W}_{j}$, $\mathcal{V}_{j}^{n}\cap\mathcal{V}_{j}^{n^{\prime}}=\emptyset$, $\forall n\neq n^{\prime}$;
  2. (ii).
  3. (iii).
    There exists $1\leq\hat{s}_{1}<\hat{e}_{1}\leq\hat{s}_{2}<\hat{e}_{2}\leq\dots\leq\hat{s}_{N}<\hat{e}_{N}\leq T$, and $\{g_{n}\}_{n=1}^{N}\in\left[1,\infty\right)^{N}$ such that for each $1\leq n\leq N$, $(\hat{s}_{n},\hat{e}_{n})$ is a $\left(2^{-(j+1)}g_{n}|\mathcal{V}_{j}^{n}|,2^{-(j+2)}|\mathcal{V}_{j}^{n}|,\frac{\log(2)}{2\log_{2}(k)}\right)$-segment with index set as $\mathcal{V}_{j}^{n}$. That is, the following hold for each $1\leq n\leq N$:
    • a.
      $\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j+2}}\leq\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{\hat{s}_{n}}\leq\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j+1}}$ ;
    • b.
      $\frac{g_{n}|\mathcal{V}_{j}^{n}|}{2^{j}}\cdot\exp\left(\frac{\log(2)}{2\log_{2}(k)}\right)\leq\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{\hat{e}_{n}}$;
    • c.
      $\sum_{i\in\mathcal{V}_{j}^{n}}w_{i}^{t}\geq\frac{|\mathcal{V}_{j}^{n}|}{2^{j+2}}$ for any $\hat{s}_{n}\leq t\leq\hat{e}_{n}$.

### v3

@@ D.3 Statement and proof of Lemma 19
Let $\overline{\sigma}$ be the permutation over $\mathcal{H}$ with $\overline{\sigma}(h^{\star})=h,\overline{\sigma}(h)=h^{\star}$ and $\overline{\sigma}(h^{\prime})=h^{\prime}$ for all $h^{\prime}\notin\{h,h^{\star}\}$. It then holds that $\overline{\sigma}^{-1}=\overline{\sigma}$.
Consider a given sequence $\{m_{i}\}_{i=1}^{k}$. Let $X(i)=\{X^{l}(i)\}_{l=1}^{m_{i}}\in\{-1,0,1\}^{kNm_{i}}$ for $i\in[k]$, and let $x(i)=\{x^{l}(i)\}_{l=1}^{m_{i}}$ be the datapoints of the first $m_{i}$ calls to $\mathsf{Query}(D_{i})$. With slight abuse of notation, we take $\sigma(x(i))=\{\sigma(x^{l}(i))\}_{l=1}^{m_{i}}$ for each $i\in[k]$. It then follows from Lemma 20 in Appendix D.4 that
$\mathbb{P}_{\mathcal{G},\mathcal{H}}\left\{h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right\}$
$=\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=\sigma^{-1}(h),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right\},$
which implies that
$\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h^{\star},\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\sigma(x(i))=X(i),\forall i\in[k]\right\}$
$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\overline{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right\}\cdot\frac{\mathbb{P}\left\{\sigma(x(i))=X(i),\forall i\in[k]\right\}}{\mathbb{P}\left\{\overline{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right\}}$
$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\overline{\sigma}\sigma(x(i))=X(i),\forall i\in[k]\right\}\cdot\frac{\mathbb{P}\left\{\sigma(x(\widetilde{i}))=X(\widetilde{i})\right\}}{\mathbb{P}\left\{\overline{\sigma}\sigma(x(\widetilde{i}))=X(\widetilde{i})\right\}}$
$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\overline{\sigma}\sigma(x(i)))=X(i),\forall i\in[k]\right\}$
$\quad\qquad\qquad\qquad\qquad\cdot\frac{\mathbb{P}\left\{\{x_{j_{h,i}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X_{j_{\sigma(h),\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h^{\star},\widetilde{i}}}^{l}({\widetilde{i}})\}_{l=1}^{m_{\widetilde{i}}}=\{X_{j_{\sigma(h^{\star}),{\widetilde{i}}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\right\}}{\mathbb{P}\left\{\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X_{j_{\sigma(h),\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X_{j_{\sigma(h^{\star}),\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\right\}}.$ (133)
Rearrange the equation to arrive at
$\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\Bigg{\{}h_{\mathrm{out}}=h^{\star},\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\sigma(x(i))=X(i),\forall i\in[k]$
|$\qquad\qquad\qquad\qquad\Big{|}\,\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\Bigg{\}}$
$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\Bigg{\{}h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k},\overline{\sigma}\sigma(x(i))=X(i),\forall i\in[k]$
|$\qquad\qquad\qquad\quad\Big{|}\,\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\Bigg{\}}.$
Taking the sum over all possible choices of $\{X(i)\}_{i\neq\widetilde{i}}$, $\left\{\{X^{l}_{j_{h^{\prime},i^{\prime}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\right\}_{i^{\prime}\in[k],h^{\prime}\notin\{h,h^{\star}\}}$,
$\left\{\{X^{l}_{j_{h^{\prime},i^{\prime}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\right\}_{h^{\prime}\in\{h,h^{\star}\},i^{\prime}\neq\widetilde{i}}$ and $\{m_{i}\}_{i\neq\widetilde{i}}$, we reach
$\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\Bigg{\{}h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}=m_{\widetilde{i}}$
|$\qquad\qquad\qquad\qquad\Big{|}\,\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\Bigg{\}}$
$=\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\Bigg{\{}h_{\mathrm{out}}=h,M_{\widetilde{i}}=m_{\widetilde{i}}$
|$\qquad\qquad\qquad\qquad\Big{|}\,\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}\Bigg{\}}$
for any $X(\widetilde{i})\in\{-1,0,1\}^{kNm_{\widetilde{i}}}$.
Fix $m_{\widetilde{i}}=m$, and choose $\{X^{l}_{j_{\sigma(h),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=v_{1}$, $\{X^{l}_{j_{\sigma(h^{\star}),\widetilde{i}}}(\widetilde{i})\}_{l=1}^{m_{\widetilde{i}}}=v_{2}$. We then have
$\mathbb{P}_{\mathcal{G}^{\prime},\mathcal{H}}\left\{h_{\mathrm{out}}=h^{\star},M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h^{\star},M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathbb{P}_{\mathcal{G},\overline{\sigma}\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=\overline{\sigma}^{-1}(h^{\star}),M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1}\right\}$
$=\frac{1}{|\Pi_{\mathcal{H}}|}\sum_{\sigma\in\Pi_{\mathcal{H}}}\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=h,M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1}\right\}$
$=\mathbb{P}_{\mathcal{G}^{\prime},\mathcal{H}}\left\{h_{\mathrm{out}}=h,M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1}\right\}.$
This completes the proof. ∎
D.4 Statement and proof of Lemma 20
@@
Consider any $\{m_{i}\}_{i\in[k]}$, $\sigma\in\Pi_{\mathcal{H}}$ and $X\in\{-1,0,1\}^{kN\sum_{i=1}^{k}m_{i}}$. Let $\{X(i)\}_{i\in[k]}$ and $\{x(i)\}$ be defined as in Lemma 19. It then holds that
$\mathbb{P}_{\mathcal{G},\mathcal{H}}\left[h_{\mathrm{out}}=h,\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid x(i)=X(i),\forall i\in[k]\right]$
$=\mathbb{P}_{\mathcal{G},\sigma(\mathcal{H})}\left\{h_{\mathrm{out}}=\sigma^{-1}(h),\{M_{i}\}_{i=1}^{k}=\{m_{i}\}_{i=1}^{k}\mid\sigma^{-1}(x(i))=X(i),\forall i\in[k]\right\}.$ (134)